OpenStudy (anonymous):

Solve for X : ((3)/(5))^(x)=7^(1-x)

OpenStudy (lgbasallote):

\[\huge \frac{3}{5^x} = 7^{1-x}\] is that the question?

OpenStudy (lgbasallote):

or is it \[\huge \left(\frac 35 \right)^x = 7^{1-x}\]

OpenStudy (anonymous):

the second one

OpenStudy (anonymous):

didnt help

OpenStudy (anonymous):

split the left side, and then take the log of both sides

OpenStudy (anonymous):

\[ln(3^x)-ln(5^x)=ln(7^{1-x})\]

OpenStudy (anonymous):

pull out the x to the front \[xln(3)-xln(5)=(1-x)ln(7)\]

OpenStudy (anonymous):

distribute the right side \[xln(3)-xln(5)=ln(7)-xln(7)

OpenStudy (anonymous):

\[xln(3)-xln(5)=ln(7)-xln(7)\]

OpenStudy (anonymous):

get all the x terms on one side and factor an x out

OpenStudy (anonymous):

can you finish this

OpenStudy (anonymous):

still stuck >.<

OpenStudy (waleed_imtiaz):

Did that help ?

OpenStudy (anonymous):

\[x \ln (3/5)=(1-x)\ln 7\] \[x \ln 3-x \ln 5=\ln 7-xln 7\] like terms \[x \ln 3-xln 5+x \ln 7=\ln 7\] \[x \ln (3*7/5)=\ln 7\] \[x=\ln (21/5)/\ln 7\]

OpenStudy (anonymous):

Alternatively \[(3/5)^x=7/7^x\] \[(3/5)^x(7^x)=7\] \[(3/5*7)^x=7\] \[x=\log_{21/7}7 \] which is also \[x \ln (3/5*7)=\ln 7\]

OpenStudy (anonymous):

That's not one of my choices. my choices are A. No real solutions B. − 5.76 C. − 0.8599 D. 1.356 E. None of the above

OpenStudy (anonymous):

(3/5)^x = 7^(1 - x) x log (3/5) = (1 - x ) log 7 x log (3/5) = log 7 - x log 7 x [ log (3/5) + log 7 ] = log 7 x [ log (21/5) ] = log 7 x = log 7 / log [ (21/5) ] x = 1.356

OpenStudy (anonymous):

thanks love.

OpenStudy (anonymous):

ur welcome

hero (hero):

I could have posted that as a solution, but I was too busy trying to challenge myself to express 3/5 as a power of seven.

OpenStudy (anonymous):

lol :P